Run Planning Meeting Notes, Feb 14 - Feb 20, 2019

From GlueXWiki
Jump to: navigation, search

Thursday, Feb. 14 at 08:45

  1. Past 24 hours:
  2. Issues
  3. Today:
    • No planned access
    • Opportunistic access requested by Kenoe to work on GEM.
    • DIRC nominal running (MOROC threshold scan?)

Friday, Feb. 15 at 08:45

  1. Past 24 hours:
    • Nominal DIRC setting collected 1B triggers
      • Switched to the higher MOROC threshold (200)
    • Smooth overnight running at 180nA, coming back to the desired 225nA this morning.
    • Both opportunistic accesses happened during the accelerator down time due to the bulk power supply failure.
  2. Issues:
  3. Plan for today:
    • keep collecting higher MOROC threshold data (up to 1B triggers).
    • When beam is not available, 15 min LED runs with different pulsing and trigger conditions.
    • No planned and opportunistic access

Monday, Feb. 18 at 08:45

  1. Over the weekend summary:
    • Quote by Eugine: "All the problems landed on other ppl's court".
    • DIRC SSP firmware on Friday (thanks to Ben and Sasha)
    • Low intensity program completed.
    • HV scan: two more settings remains
    • HI test on Sunday: for most of the day, we were the only hall up.
      • ToF inner layer off, DC HV reduced, Tpol excluded from the trigger.
      • DIRC and DAQ passed the test!
    • Overnight: smooth 425nA scanning operation.
  2. Monitoring
    • Only 6 out of 30 root monitoring plots are fully occupied.
  3. Plan for today:
    • Hall A, B and C are down, another HI test? @900nA?
  4. Issues:
    • hdview2: hang on the 1st event
    • DAQ: Randy factor cant be set and BCALroc.
    • Any update on the RF issue?
  5. Transition to the CPP

Tuesday, Feb. 19 at 08:45

  1. Past 24 hours
    • 500 nA threshold scan: trigger rate 75k Hz
    • Overnight: smooth 350nA scanning operation.
    • HV scan: almost done, only few more settings remains
  2. Data Analysis:
  3. Monitoring and rootspy:
  4. Plan for today:
    • Complete HV scan and finish 350nA threshold scan.
  5. Issues:
  6. Hall access today:
    • RF recovery for 8 hours.
    • Mark, Colin and Beni. Restricted or not?
    • Tagger Hall: restricted by 10 am for neutron detector work
  7. Transition to the CPP
    • Hall Access starting 7am tomorrow (Feb/19).
    • Any running plan? As a matter of fact, yes.
    • No RC meeting tomorrow.

Wednesday, Feb. 20 at 08:45

  1. DIRC commissioning completed, and DIRC crews are happy.
  2. Issues:
  3. CPP tasks and schedule:
    • Revert back from the DIRC lumi settings (FDC, ToF, TAGM, TPol).
    • Currently under restricted access til 12:00 pm, hall work on-going
    • Lockup at noon (11:45am, MCC announcement).
    • Radcon support arranged for 3-6pm
  4. CPP run plan:
  5. MCC whiteboard: